OpenStudy (freemap):

Please Help Me 8b – 1 = 24b + 4 b =

OpenStudy (mathstudent55):

You have the variable b on both sides. You also have numbers, -1 and 4, on both sides.

OpenStudy (kendricklamar2014):

Step 1. Subtract 24 from both sides: 8b−1−24b=24b+4−24b = −16b−1=4

OpenStudy (michele_laino):

If I subtract \(4\) from both sides, I get: \[8b - 1 - 4 = 24b + 4 - 4\] please simplify

OpenStudy (mathstudent55):

You need all variables on one side, and all numbers on the other side.

OpenStudy (kendricklamar2014):

Step 2. Add 1 to both sides: −16b−1+1=4+1 −16b=5

OpenStudy (michele_laino):

since \(-1-4=-5\) and \(4-4=0\) then i can write: \[8b - 5 = 24b\]

OpenStudy (idealist10):

8b-1=24b+4 8b-24b-1=4 -16b-1=4 -16b=4+1 -16b=5 b=5/-16 b=-5/16 Answer: -5/16

OpenStudy (freemap):

-0.3125

OpenStudy (michele_laino):

correct! @freemap it is \(b=-5/16=-0.3125\)

OpenStudy (idealist10):

That's why, -5/16=-0.3125, they're the same.

OpenStudy (freemap):

Thanks to everyone that helped towards this Question

OpenStudy (mathstudent55):

\(8b-1 = 24b + 4\) Step 1. Add 1 to both sides: \(8b - 1 \color{red}{+ 1} = 24b + 4 \color{red}{+ 1}\) \(8b = 24b + 5\) Step 2. subtract 24b from both sides: \(8b \color{red}{- 24b} = 24b \color{red}{- 24b} + 5\) \(-16b = 5\) Step 3. Divide both sides by -16: \(\dfrac{-16b}{\color{red}{-16}} = \dfrac{5}{\color{red}{-16}} \) \(b = -\dfrac{5}{16} \)

OpenStudy (freemap):

Thanks for the illustration

OpenStudy (idealist10):

By the way, can anyone help me with a differential equations problem? http://openstudy.com/study#/updates/562540b9e4b042c91fffc8a3

OpenStudy (freemap):

I would but I'm not good at math, so sorry.